OpenStudy (anonymous):

it applies for 95 also

OpenStudy (anonymous):

is this a question?

OpenStudy (anonymous):

9^2=81 +1=91 ADD 25=>9125 BUT95^2=9025

OpenStudy (anonymous):

5*5=25 9*10=90 so answer is 9025

OpenStudy (anonymous):

ok,i misunderstood the idea:)

OpenStudy (mathmate):

It works for any number finishing in a five. n|5^2=(n)(n+1)|25 where | stands for a concatenation. 125^2=12*13|25=15625 i.e. 12*13 =156, tag on 25 at the end to give 15625.